# On the Curses of Future and History in Future-dependent Value Functions for Off-policy Evaluation

### v1

@@ 5 Effective History Weights and A New Algorithm
We now turn to Q2 in Section 3, which asks for a quantitative understanding of the $\textrm{IV}(\mathcal{V})$ term. Note that the $\textrm{IV}(\mathcal{V})$ and the $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ terms, taken together, are to address the conversion from what we can estimate and minimize:
$\sqrt{\mathbb{E}_{\pi_{b}}[(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})^{2}]},$ (8)
to what we want to bound (see Lemma 1), which differs both in the policy ($\pi_{b}\to\pi_{e}$) and the operator ($\mathcal{B}^{\mathcal{H}}\to\mathcal{B}^{{\mathcal{S}}}$):
$\mathbb{E}_{\pi_{e}}[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})].$ (9)
While we could directly define a parameter by taking the worst-case (over $V\in\mathcal{V}$) ratio between the two expressions, [^7] Recent offline RL theory indeed employed such definitions (Xie et al., 2021; Song et al., 2022), but that is after we developed mature understanding of their properties, such as being upper bounded by density ratios. the real question is to provide more intuitive understanding of when it can be bounded.
Towards this goal, Uehara et al. (2022) split the above ratio into two terms, $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$, which take care of the $\mathcal{B}^{\mathcal{H}}\to\mathcal{B}^{{\mathcal{S}}}$ conversion (under $\pi_{b}$) and $\pi_{b}\to\pi_{e}$ conversion (under $\mathcal{B}^{{\mathcal{S}}}$), respectively. While this leads to an intuitive upper bound of the $\pi_{b}\to\pi_{e}$ conversion parameter in terms of latent state coverage, the nature of the $\mathcal{B}^{\mathcal{H}}\to\mathcal{B}^{{\mathcal{S}}}$ conversion term, $\textrm{IV}(\mathcal{V})$, remains mysterious.
In this section, we take a different approach by directly providing an intuitive upper bound on both conversions altogether, under a novel belief coverage assumption. We will also revisit the split into $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$: in the absence of strong structures from $\mathcal{V}$, the boundedness of $\textrm{IV}(\mathcal{V})$ turns out to require an even stronger version of belief coverage, rendering the split unnecessary. In addition to providing an intuitive understanding to the conversion ratio, our approach also leads to a novel algorithm for estimating $J(\pi_{e})$ that replaces Bellman-completeness (Assumption 6) with a weight-realizability assumption, similar to MIS estimators for MDPs (Liu et al., 2018; Uehara et al., 2020).
@@ 5.1 Effective History Weights
The effective history weights $w^{\star}:\mathcal{H}\to\mathbb{R}$ is any function that satisfies the following for all $V\in\mathcal{V}$ and $h\in[H]$:
@@
We first see that well-bounded $w^{\star}$ immediately leads to a good conversion ratio between Eqs.(8) and (9):
$~{}\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})\right]=\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})]$
$\leq$ $~{}\sqrt{\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})^{2}]\mathbb{E}_{\pi_{b}}[(\mathcal{B}^{\mathcal{H}}V)(\tau)^{2}]},$
where the second step follows from Cauchy-Schwartz for r.v.’s. As can be seen, all we need is $\|w^{\star}\|_{2,d^{\pi_{b}}_{h}}:=\sqrt{\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})^{2}]}$, the $\pi_{b}$-weighted 2-norm of $w^{\star}$, to be bounded, and we focus on this quantity next.
Similar to Section 4, there may be multiple $w^{\star}$ that satisfies the definition, and we only need to show the boundeness of any solution. The easiest way to bound $w^{\star}$ is to provide an explicit construction. Also similarly, the most obvious construction is $w^{\star}(\tau_{h})=\frac{\Pr_{\pi_{e}}(\tau_{h})}{\Pr_{\pi_{b}}(\tau_{h})}=\prod_{h^{\prime}=1}^{h-1}\frac{\pi_{e}(a_{h^{\prime}}|o_{h^{\prime}})}{\pi_{b}(a_{h^{\prime}}|o_{h^{\prime}})}$, noting that $\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})\right]=\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})\right]$ and importance weighting on $\tau_{h}$ changes the distribution from $\mathbb{E}_{\pi_{b}}$ to $\mathbb{E}_{\pi_{e}}$. However, the use of cumulative importance weights is undesirable given its exponential nature, causing the history-version of “curse of horizon” (Liu et al., 2018).
@@
We now show a better construction that is bounded under a natural belief coverage assumption. Note that Eq.(10) can be written as:
@@
where $\mathcal{B}^{{\mathcal{S}}}_{h}V\in\mathbb{R}^{S}$ is the Bellman residual vector for $V$ on ${\mathcal{S}}_{h}$. As a sufficient condition (which is also necessary when $\mathcal{V}$ lacks strong structures, i.e., $\{\mathcal{B}^{{\mathcal{S}}}_{h}V:V\in\mathcal{V}\}$ spans the entire $\mathbb{R}^{S}$), we can find $w^{\star}$ that satisfies:
$\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})\mathbf{b}(\tau_{h})]=\mathbb{E}_{\pi_{e}}\left[\mathbf{b}(\tau_{h})\right]=:\mathbf{b}_{h}^{\pi_{e}}.$
This is related to the mean matching problem in the distribution shift literature (Gretton et al., 2009; Yu & Szepesvári, 2012), and a linear-regression-based solution is (Bruns-Smith et al., 2023):
@@
where $\Sigma_{\mathcal{H},h}:=\sum_{\tau_{h}}d^{\pi_{b}}(\tau_{h})\mathbf{b}(\tau_{h})\mathbf{b}(\tau_{h})^{\top}$, and $\mathbf{b}_{h}^{\pi_{e}}$ coincides with the latent state distribution $[d^{\pi_{b}}(s_{h})]_{s_{h}}$.
@@
Under 10, $w^{\star}$ in Eq. (11) satisfies $\forall h$, $\|w^{\star}\|^{2}_{2,d^{\pi_{b}}_{h}}\leq C_{\mathcal{H},2}.$
See Appendix D.1 for the proof. 10 requires that the covariance matrix of belief states under $\pi_{b}$ covers $\mathbf{b}^{\pi_{e}}$, the average belief state under $\pi_{e}$. As before, we also check the on-policy case (see Appendix D.2):
@@
As already discussed below Proposition 4 in Section 4, this is also analogous to the feature coverage assumption in OPE for linear MDPs.
Comparison to $\textrm{IV}(\mathcal{V})$
We now go back to the $\textrm{IV}(\mathcal{V})$ term and connect it to our assumption. Perhaps surprisingly, we show in the proposition below that, under moderate assumptions, $\textrm{IV}(\mathcal{V})$ must scale with $\frac{1}{\sigma_{\min}(\Sigma_{\mathcal{H},h})}$.
Proposition 6.
The proof is deferred to Appendix D.3. The proposition states that as long as $\min_{s_{h}}d^{\pi_{b}}(s_{h})$ is bounded away from 00 (which is benign and helps bound the $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ term) and $\mathcal{V}$ is sufficiently rich that $\mathcal{B}^{{\mathcal{S}}}_{h}\mathcal{V}$ includes a certain vector in $\mathbb{R}^{S}$, then boundedness of $\textrm{IV}(\mathcal{V})$ requires bounded $1/\sigma_{\min}(\Sigma_{\mathcal{H},h})$, which is stronger than our belief coverage assumption. Therefore, the split between $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$ is superficial.
5.2 $L_{\infty}$ Coverage Assumptions
While we have all the assumptions to overcome the curses of future (Section 4) and history (Section 5.1), the outcome coverage $C_{\mathcal{F},2}$ incurs $S$ dependence (Example 3), and the boundedness of FDVF requires an additional regularity assumption (8). In fact, it will also be nice to obtain $\|w^{\star}\|_{\infty}$ boundedness, as we will rely on it in the new algorithm in Section 5.3; however, Assumption 10 only guarantees the boundedness of $\|w^{\star}\|_{2,d^{\pi_{b}}_{h}}^{2}$ (Lemma 5), unless we make yet another regularity assumption on belief vectors.
Motivated by these considerations, we introduce the $L_{\infty}$ versions of the coverage assumptions below, which help bound $\|V_{\mathcal{F}}\|_{\infty}$ and $\|w^{\star}\|_{\infty}$ without $S$ scaling or additional regularity conditions. We will also compare them to the $L_{2}$ versions and discuss which version is more appropriate depending on the use cases.
To develop intuitions, we first examine the boundedness of $V_{\mathcal{F}}$ in the setting of Example 2. In this example, $\Sigma_{\mathcal{F},h}=\mathbf{I}$, and $\|V_{\mathcal{F}}\|_{\infty}\leq H$. According to the $L_{2}$ Hölder decomposition in Proposition 4, however,
The comparison clearly highlights that the looseness in $S$ comes from the fact that $L_{2}$ Hölder does not fully leverage the fact that $\|\mathbf{u}(f_{h})/Z(f_{h})\|_{1}=1$ and loosely relaxing it to $\|\mathbf{u}(f_{h})/Z(f_{h})\|_{2}\leq 1$.
$L_{\infty}$ Belief Coverage
The above example suggests that we may want to consider $L_{1}/L_{\infty}$ Hölder for $L_{\infty}$ boundedness. We start with belief coverage which is simpler. To bound $\|w^{\star}\|_{\infty}$, we propose the following decomposition:
$|w^{\star}(\tau_{h})|\leq\|\mathbf{b}(\tau_{h})\|_{1}\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}=\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}.$
This way, we can immediately bound $\|w^{\star}\|_{2,d^{\pi_{b}}_{h}}^{2}\leq\|w^{\star}\|_{\infty}$ with the following assumption:
@@ 5.2 $L_{\infty}$ Coverage Assumptions
Assume $\forall h$,
$\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}\leq C_{\mathcal{H},\infty}.$
As a direct consequence, $\|w^{\star}\|_{\infty}\leq C_{\mathcal{H},\infty}$.
A question the readers might have is why the decomposition $\|\mathbf{b}(\tau_{h})\|_{1}\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}$ is any better than other arbitrary splits. In particular, $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$ is the inverse of second moment (covariance) multiplying the first moment (expectation), raising the concern that the quantity may be poorly scaled: for example, given bounded (and otherwise unconstrained) random vector $x$, $\mathbb{E}[xx^{\top}]^{-1}\mathbb{E}[x]$ can go to infinity if we rescale $x$ by a small constant close to 00.
We first observe that such a pathology will not happen with $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$ because the random vectors ($\mathbf{b}(\tau_{h})$) are $L_{1}$-normalized, which is exactly the property we seek to leverage. Below we use a few examples to show that $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$ is a very well-behaved quantity, starting with the on-policy case. Perhaps surprisingly, $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$ has an exact solution in this situation:
@@
When $\pi_{e}=\pi_{b}$, $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}=\mathbf{1}$, the all-one vector; see Appendix D.4 for the calculation. Consequently, Assumption 11 is satisfied with $C_{\mathcal{H},\infty}=1$.
It is also instructive to consider the case when $\mathbf{b}(\tau_{h})$ is always one-hot, i.e., histories always reveal the latent state. (This is analogous to Example 2.)
@@
In this special case, $\Sigma_{\mathcal{H},h}$ is a diagonal matrix where the diagonal is the latent state distribution induced by $\pi_{b}$, whereas $\mathbf{b}_{h}^{\pi_{e}}$ is that for $\pi_{e}$. In this case, $\|\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}$ is a standard measure of coverage using the infinity-norm of the density ratio (Munos, 2007; Chen & Jiang, 2019; Uehara et al., 2022). In comparison, we can also calculate $\|\mathbf{b}^{\pi_{e}}\|^{2}_{\Sigma_{\mathcal{H},h}^{-1}}$ from Assumption 10, which turns out to be $\mathbb{E}_{\pi_{b}}[(d^{\pi_{e}}(s_{h})/d^{\pi_{b}}(s_{h}))^{2}]$. The 2nd moment of the density ratio is also a measure of coverage, which is tighter than the $L_{\infty}$ measure (Xie & Jiang, 2020). In fact, this is generally true even when $\mathbf{b}(\tau_{h})$ is not one-hot: (see proof in Appendix D.5)
@@
$\|\mathbf{b}_{h}^{\pi_{e}}\|_{\Sigma_{\mathcal{H},h}^{-1}}^{2}\leq\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}.$
Therefore, if we only need $\|w^{\star}\|_{2,d^{\pi_{b}}_{h}}^{2}$ boundedness, the $L_{2}$ belief coverage (10) is generally a tighter assumption, and the $L_{\infty}$ belief coverage above (11) is useful when we want bounded $\|w^{\star}\|_{\infty}$.
$L_{\infty}$ Outcome Coverage
We now show how to improve the analysis on the boundedness of $V_{\mathcal{F}}$ from Section 4 in a similar manner. To leverage the lesson we have learned on belief coverage, we would like to also express the coverage assumption for $V_{\mathcal{F}}$ in terms of covariance and expectation of stochastic vectors. Luckily, the construction in Eq.(7) is already in this form, as $Z_{h}^{-1}M_{\mathcal{F},h}^{\top}$ has rows that sum up to $1$. Let $\bar{\mathbf{u}}(f_{h}):=\mathbf{u}(f_{h})/Z(f_{h})\in\Delta({\mathcal{S}}_{h})$, then $\Sigma_{\mathcal{F},h}=S\cdot\mathbb{E}_{Z_{h}/S}[\bar{\mathbf{u}}\bar{\mathbf{u}}^{\top}]$. While we could also do a similar $L_{1}/L_{\infty}$ decomposition, it is difficult to bound $\Sigma_{\mathcal{F},h}^{-1}V_{{\mathcal{S}}}^{\pi_{e}}$ in the on-policy case. Compared to the situation of $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$, we see the difference is that both $\Sigma_{\mathcal{H},h}^{-1}$ and $\mathbf{b}_{h}^{\pi_{e}}$ are properties of $\mathbf{b}(\tau_{h}),\tau_{h}\sim\pi_{b}$; in contrast, $V_{{\mathcal{S}}}^{\pi_{e}}$ depends on the reward function but $\Sigma_{\mathcal{F},h}$ is ignorant of it, and this mismatch makes the on-policy boundedness difficult.
Based on this observation, we propose to modify the construction in Eq.(7) slightly to incorporate reward information: define $Z^{R}(f_{h}):=Z(f_{h})/R(f_{h})$, and we use $Z^{R}$ to replace $Z$ in Eq.(7):
$V_{\mathcal{F},h}=(Z_{h}^{R})^{-1}M_{\mathcal{F},h}^{\top}(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}},h}^{\pi_{e}},$ (12)
$\textrm{where~{}}\Sigma_{\mathcal{F},h}^{R}:=M_{\mathcal{F},h}(Z_{h}^{R})^{-1}M_{\mathcal{F},h}^{\top}.$
We now perform the $L_{1}/L_{\infty}$ decomposition similar to the case of history, and show that the resulting coverage condition satisfies desirable properties.
Assumption 12 ($L_{\infty}$ outcome coverage).
Assume for all $h$,
$\|(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}\|_{\infty}\leq C_{\mathcal{F},\infty}.$
Lemma 8.
Under 12, $\|V_{\mathcal{F}}\|_{\infty}\leq HC_{\mathcal{F},\infty}$.
See proof in Appendix D.6. We first examine the on-policy case, which closely resembles Example 6 and the construction exactly recovers $V_{\mathcal{F}}=R$, a property that the construction in Eq.(7) does not enjoy; see Appendix D.7 for details.
Example 8.
When $\pi_{e}=\pi_{b}$, $(\Sigma_{\mathcal{F},h}^{R})^{-1}V_{{\mathcal{S}}}^{\pi_{e}}=\mathbf{1}$, thus 12 holds with $C_{\mathcal{F},\infty}=1$. Furthermore, the construction in Eq.(12) is exactly $V_{\mathcal{F}}=R$.
Next we examine the setting of Example 2; see the calculation in Appendix D.8.
Example 9.
[…]
While $(\Sigma_{\mathcal{F},h}^{R})^{-1}$ may behave poorly if $V_{{\mathcal{S}}}^{\pi_{b}}(s_{h})$ is small, an easy fix is to simply add a constant to the reward function and shift its range to e.g., [1, 2], which ensures a small $C_{\mathcal{F},\infty}$.
Algorithm Guarantee
Now we have all the pieces to present a fully polynomial version of Theorem 2 under the proposed coverage assumptions. One subtlety is that the guarantee depends on $C_{\mathcal{V}}:=\max_{V\in\mathcal{V}}\|V\|_{\infty}$, which is closely related to $\|V_{\mathcal{F}}\|_{\infty}$ since we require $V_{\mathcal{F}}\in\mathcal{V}$, but they are not equal since $\mathcal{V}$ can include other functions with higher range. To highlight the dependence of $\|V_{\mathcal{F}}\|_{\infty}$ on the proposed coverage assumptions, we follow Xie & Jiang (2020) to assume that the range of the function classes is not much larger than that of the function it needs to capture. The proof of the theorem is deferred to Appendix D.9.
Theorem 9.
Consider the same setting as Theorem 2, and let Assumptions 10 and 12 hold. For some absolute constant $c$, further assume $C_{\mathcal{V}}\leq c\|V_{\mathcal{F}}\|_{\infty}$ for $V_{\mathcal{F}}$ in Eq.(12) and $C_{\Xi}\leq c(\|V_{\mathcal{F}}\|_{\infty}+1)$. W.p. $\geq 1-\delta$,
$~{}|J(\pi_{e})-\mathbb{E}_{\mathcal{D}}[\widehat{V}(f_{1})]|$
$\leq$ $~{}cH^{2}(C_{\mathcal{F},\infty}+1)\sqrt{\frac{C_{\mathcal{H},2}C_{\mu}\log(|\mathcal{V}||\Xi|/\delta)}{n}}.$
In addition to $H$ and complexities of $|\mathcal{V}|$ and $|\Xi|$, the bound only depends on the intuitive coverage parameters: $C_{\mu}$ (action coverage), $C_{\mathcal{H},2}$ ($L_{2}$ belief coverage, which is tighter than the $L_{\infty}$ version and we do not need bounded $\|w^{\star}\|_{\infty}$ here), and $C_{\mathcal{F},\infty}$ ($L_{\infty}$ outcome coverage).

### v2

@@ 5 Effective History Weights and A New Algorithm
We now turn to Q2 in Section 3, which asks for a quantitative understanding of the $\textrm{IV}(\mathcal{V})$ term. Note that $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$, taken together, are to address the conversion between:
[…]
The two terms differ both in the policy ($\pi_{b}\to\pi_{e}$) and the operator ($\mathcal{B}^{\mathcal{H}}\to\mathcal{B}^{{\mathcal{S}}}$). While we could directly define a parameter by taking the worst-case (over $V\in\mathcal{V}$) ratio between the two expressions, [^5] Recent offline RL theory indeed employed such definitions (Xie et al., 2021; Song et al., 2022), but that is after we developed mature understanding of their properties, such as being upper bounded by density ratios. the real question is to provide more intuitive understanding of when it can be bounded.
Towards this goal, Uehara et al. (2022a) split the above ratio into two terms, $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$, which take care of the $\mathcal{B}^{\mathcal{H}}\to\mathcal{B}^{{\mathcal{S}}}$ conversion (under $\pi_{b}$) and $\pi_{b}\to\pi_{e}$ conversion (under $\mathcal{B}^{{\mathcal{S}}}$), respectively. While this leads to an intuitive upper bound of the $\pi_{b}\to\pi_{e}$ conversion parameter in terms of latent state coverage, the nature of the $\mathcal{B}^{\mathcal{H}}\to\mathcal{B}^{{\mathcal{S}}}$ conversion, $\textrm{IV}(\mathcal{V})$, remains mysterious.
In this section, we take a different approach by directly providing an intuitive upper bound on both conversions altogether, under a novel belief coverage assumption. In Appendix E.4 we will also revisit the split into $\textrm{IV}(\mathcal{V})$ and $\textrm{Dr}_{\mathcal{V}}[d^{\pi_{e}},d^{\pi_{b}}]$: in the absence of strong structures from $\mathcal{V}$, the boundedness of $\textrm{IV}(\mathcal{V})$ turns out to require an even stronger version of belief coverage, rendering the split unnecessary. Furthermore, our approach also leads to a novel algorithm for estimating $J(\pi_{e})$ that replaces Bellman-completeness (Assumption 6) with a weight-realizability assumption, similar to MIS estimators for MDPs (Liu et al., 2018; Uehara et al., 2020).
@@ 5.1 Effective History Weights
An effective history weight function $w^{\star}:\mathcal{H}\to\mathbb{R}$ is any function that satisfies: $\forall V\in\mathcal{V}$, $h\in[H]$,
@@
We first see that well-bounded $w^{\star}$ immediately leads to a good conversion ratio for Eq.(6):
$|\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})\right]|=|\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})]|\leq\sqrt{\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})^{2}]\mathbb{E}_{\pi_{b}}[(\mathcal{B}^{\mathcal{H}}V)(\tau)^{2}]},$
where the inequality follows from Cauchy-Schwartz for r.v.’s. Hence, all we need is $\|w^{\star}\|_{2,d^{\pi_{b}}_{h}}:=\sqrt{\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})^{2}]}$, the $\pi_{b}$-weighted 2-norm of $w^{\star}$, to be bounded, and we focus on this quantity next.
Similar to Section 4, there may be multiple $w^{\star}$ that satisfies the definition, and we only need to show the boundeness of any solution. Also similarly, the most obvious solution is $w^{\star}(\tau_{h})=\frac{\Pr_{\pi_{e}}(\tau_{h})}{\Pr_{\pi_{b}}(\tau_{h})}=\prod_{h^{\prime}=1}^{h-1}\frac{\pi_{e}(a_{h^{\prime}}|o_{h^{\prime}})}{\pi_{b}(a_{h^{\prime}}|o_{h^{\prime}})}$, noting that $\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})\right]=\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{\mathcal{H}}V)(\tau_{h})\right]$ and importance weighting on $\tau_{h}$ changes $\mathbb{E}_{\pi_{b}}$ to $\mathbb{E}_{\pi_{e}}$. However, the use of cumulative importance weights is undesirable given its exponential nature, causing the history-version of “curse of horizon” Liu et al. (2018).
@@
We now show a better construction that is bounded under a natural belief coverage assumption. Note that Def 10 can be written as:
@@
where $\mathcal{B}^{{\mathcal{S}}}_{h}V\in\mathbb{R}^{S}$ is the Bellman residual vector for $V$ on ${\mathcal{S}}_{h}$. As a sufficient condition (which is also necessary when $\mathcal{V}$ lacks strong structures, i.e., $\{\mathcal{B}^{{\mathcal{S}}}_{h}V:V\in\mathcal{V}\}$ spans the entire $\mathbb{R}^{S}$), we can find $w^{\star}$ that satisfies: $\mathbb{E}_{\pi_{b}}[w^{\star}(\tau_{h})\mathbf{b}(\tau_{h})]=\mathbb{E}_{\pi_{e}}\left[\mathbf{b}(\tau_{h})\right]=:\mathbf{b}_{h}^{\pi_{e}}.$ This is related to the mean matching problem in the distribution shift literature (Gretton et al., 2009; Yu and Szepesvári, 2012), and a standard solution is (Bruns-Smith et al., 2023):
@@
where $\Sigma_{\mathcal{H},h}:=\sum_{\tau_{h}}d^{\pi_{b}}(\tau_{h})\mathbf{b}(\tau_{h})\mathbf{b}(\tau_{h})^{\top}$, and $\mathbf{b}_{h}^{\pi_{e}}$ coincides with $[d^{\pi_{e}}(s_{h})]_{s_{h}}$.
@@
Under 11, $w^{\star}$ in Eq. (8) satisfies $\forall h$, $\|w^{\star}\|^{2}_{2,d^{\pi_{b}}_{h}}\leq C_{\mathcal{H},2}.$ See Appendix E.1.
11 requires that the covariance matrix of belief states under $\pi_{b}$ covers $\mathbf{b}^{\pi_{e}}$, the average belief state under $\pi_{e}$. As before, we also check the on-policy case (see Appendix E.2):
@@
Algorithm Guarantee
Now we have all the pieces to present a fully polynomial version of Theorem 2 under the proposed coverage assumptions. One subtlety is that the guarantee depends on $C_{\mathcal{V}}:=\max_{V\in\mathcal{V}}\|V\|_{\infty}$, which is closely related to $\|V_{\mathcal{F}}\|_{\infty}$ since we require $V_{\mathcal{F}}\in\mathcal{V}$, but they are not equal since $\mathcal{V}$ can include other functions with higher range. To highlight the dependence of $\|V_{\mathcal{F}}\|_{\infty}$ on the proposed coverage assumptions, we follow Xie and Jiang (2020) to assume that the range of the function classes is not much larger than that of the function it needs to capture. A similar assumption applies for bounding $C_{\Xi}$. The proof of the theorem is deferred to Appendix E.10.
Theorem 7.
Consider the same setting as Theorem 2, and let Assumptions 11 and 9 hold. For some absolute constant $c$, further assume $C_{\mathcal{V}}\leq c\|V_{\mathcal{F}}\|_{\infty}$ for $V_{\mathcal{F}}$ in Eq.(5) and $C_{\Xi}\leq c(\|V_{\mathcal{F}}\|_{\infty}+1)$. W.p. $\geq 1-\delta$, $|J(\pi_{e})-\mathbb{E}_{\mathcal{D}}[\widehat{V}(f_{1})]|\leq cH^{2}(C_{\mathcal{F},\infty}+1)\sqrt{\frac{C_{\mathcal{H},2}C_{\mu}\log(|\mathcal{V}||\Xi|/\delta)}{n}}$.
In addition to $H$ and complexities of $|\mathcal{V}|$ and $|\Xi|$, the bound only depends on the intuitive coverage parameters: $C_{\mu}$ (action coverage), $C_{\mathcal{H},2}$ ($L_{2}$ belief coverage), and $C_{\mathcal{F},\infty}$ ($L_{\infty}$ outcome coverage).
5.2 $L_{\infty}$ Belief Coverage
Assumption 11 enables bounded second moment of $w^{\star}$ (Lemma 6) but does not control $\|w^{\star}\|_{\infty}$, which we show will be useful for the new algorithm in Section 5.3. Here we present the $L_{\infty}$ version of belief coverage that controls $\|w^{\star}\|_{\infty}$, which also helps understand $L_{\infty}$ outcome coverage given the symmetry between history and future. As alluded to in Section 4.2 and Appendix B.5, $L_{2}$ Hölder is inappropriate for controlling the infinity-norm since it does not leverage the $L_{1}$ normalization of vectors in POMDPs. Instead, we propose the following decomposition based on $L_{1}/L_{\infty}$ Hölder: $|w^{\star}(\tau_{h})|\leq\|\mathbf{b}(\tau_{h})\|_{1}\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}=\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}.$ This way, we can immediately bound $\|w^{\star}\|_{\infty}$ with the following assumption:
@@ 5.2 $L_{\infty}$ Belief Coverage
Assume $\forall h$, $\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}\leq C_{\mathcal{H},\infty}.$ Then $\|w^{\star}\|_{\infty}\leq C_{\mathcal{H},\infty}$.
$\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$ is the inverse of second moment (covariance) multiplying the first moment (expectation), raising the concern that the quantity may be poorly scaled: for example, given bounded (but otherwise arbitrary) random vector $X$, $\mathbb{E}[XX^{\top}]^{-1}\mathbb{E}[X]$ can go to infinity if we rescale $X$ by a small constant. First note that such a pathology cannot happen here because the random vectors ($\mathbf{b}(\tau_{h})$) are $L_{1}$-normalized and cannot be arbitrarily rescaled. Below we use a few examples to show that $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$ is a very well-behaved quantity, and naturally generalize familiar concepts such as concentrability coefficient from MDPs (Munos, 2007; Chen and Jiang, 2019; Uehara et al., 2022a).
We start by checking the on-policy case. Perhaps surprisingly, $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}$ has an exact solution:
@@
When $\pi_{e}=\pi_{b}$, $\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}=\mathbf{1}$, the all-one vector; see Appendix E.5 for the calculation. Consequently, Assumption 12 is satisfied with $C_{\mathcal{H},\infty}=1$.
The next scenario considers when $\mathbf{b}(\tau_{h})$ is always one-hot, i.e., histories reveal the latent state (this is analogous to Example 2. $L_{\infty}$ coverage reduces to the familiar concentrability coefficient, the infinity-norm of density ratio as a standard coverage parameter:
@@
We can also calculate $\|\mathbf{b}^{\pi_{e}}\|^{2}_{\Sigma_{\mathcal{H},h}^{-1}}$ from Assumption 11, which equals $\mathbb{E}_{\pi_{b}}[(d^{\pi_{e}}(s_{h})/d^{\pi_{b}}(s_{h}))^{2}]$ in this “1-hot belief” scenario. Xie and Jiang (2020) show that this is tighter than $\max_{s_{h}}d^{\pi_{e}}(s_{h})/d^{\pi_{b}}(s_{h})$, and this relation extends elegantly to general belief vectors in our setting:
@@
$\|\mathbf{b}_{h}^{\pi_{e}}\|_{\Sigma_{\mathcal{H},h}^{-1}}^{2}\leq\|\Sigma_{\mathcal{H},h}^{-1}\mathbf{b}_{h}^{\pi_{e}}\|_{\infty}.$ See proof in Appendix E.6.